A truck has a 1.6 ton weight capacity in its bed. Approximately how many 2 kilogram party toys can the truck carry? (Assume volume is not a constraint and do not include units in your answer. Round to the nearest whole number.) Note: There are 0.454 kilograms in a pound and 2,000 pounds in a ton.
Answer #1

Since 1 Ton = 2000 pounts

So, 1.6 Ton = 1.6*2000 pounds = 3200 pounds

Capacity of Truck = 1.6 Ton = 3200 pounds

0.454 kilograms = 1 pound

2 Kilograms = 2/0.454 pounds = 4.40528634 pounds

Weight of one part toy = 2 Kilograms = 4.40528634 pounds

No. of Toys that can fitted in truck = Truck Capacity/Weight of one toy

No. of Toys that can fitted in truck = 3200/4.40528634

No. of Toys that can fitted in truck = 726.4

Since, no. of toys cannnot be in decimal places.

The truck can carry approximately 726 two kilogram party toys